# Generalized Finite-time Optimal Control Framework in Stochastic Thermodynamics

### v2

@@ Abstract
Optimal processes in stochastic thermodynamics are a frontier for understanding the control and design of non-equilibrium systems, with broad practical applications in biology, chemistry, and nanoscale/mesoscale systems. Optimal mass transport theory and thermodynamic geometry have emerged as optimal control methodology, but they are based on slow-driving and close to equilibrium assumptions. An optimal control framework in stochastic thermodynamics for finite time driving is still elusive. Therefore, we solve in this paper an optimal control problem for changing the control parameters of a discrete-state far-from-equilibrium process from an initial to a final value in finite-time. Optimal driving protocols are derived that minimize the total finite-time dissipation cost for the driving process. Our framework reveals that discontinuous endpoint jumps are a generic, model-independent physical mechanism that minimizes the optimal driving entropy production, whose importance is further amplified for far-from-equilibrium systems. The thermodynamic and dynamic physical interpretation and understanding of discontinuous endpoint jumps is formulated. An exact mapping between the finite-time to slow driving optimal control formulation is elucidated, developing the state-of-the-art of optimal mass transport theory and thermodynamic geometry, which has been the current paradigm for studying optimal processes in stochastic thermodynamics that relies on slow driving assumptions. Our framework opens up a plethora of applications to the thermodynamically efficient control of a far-from-equilibrium system in finite-time, which opens up a way to their efficient design principles.
@@ 1 Introduction
The framework of Stochastic Thermodynamics (ST) has developed the thermodynamic understanding of mesoscopic systems [1, 2, 3, 4, 5]. In ST, entropy production rate (EPR) quantifies the thermodynamic dissipation cost and constrains the dynamics of non-equilibrium systems. For example, the fluctuation relation (FR) and the thermodynamic kinetic uncertainty relation (TKUR) have revealed fundamental thermodynamic laws, valid beyond the second law of thermodynamics [1, 6]. Recently, a new paradigm of optimal control problems has emerged in ST. Here, the optimal control problem is loosely defined as changing an initial state (control parameter) to a final state (control parameter) in a finite time, to compute the optimal driving protocols (change of state/control parameters) that minimize the thermodynamic dissipation. Three classes of methodologies have emerged to study optimal control.
@@
First, optimal processes that minimize finite-time dissipation have been investigated for a stochastic particle in a harmonic trap [7]. Surprisingly, the finite-time optimal protocols were found to exhibit discontinuities at the initial and final times, namely the ‘kinks’. Recently, this mechanism has been understood as a mathematical artefact of the imposed boundary condition [8]. However, these works fail to capture the physical and thermodynamic origin of the ‘kinks[7, 8]. The universality of ‘kinks’ have been observed in some computationally solvable models [9, 10, 11, 12]. Despite a few model-specific studies, analytically computable solutions for other models are lacking, which has created a void in the theoretical physical understanding/framework to attribute this phenomenon.
@@
Second, thermodynamic length defines the distance in the control parameter space of a model and connects it to the thermodynamic dissipation cost [13, 14, 15, 16]. This Riemannian geometric structure in the control parameter space has been exploited to compute optimal driving protocols and formulate the framework of thermodynamic geometry, which is valid in the slow-driving limit. Within this framework, a geodesic is a minimum-distance path between the initial and final control parameters and is equivalently the optimal driving protocol [16, 17]. Thermodynamic geometry has advantages because of its practical applicability. In particular, metric tensors are numerically/experimentally computed for sophisticated models and systems [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 27]. Therefore, because of its practical applicability, thermodynamic geometry has been rigorously studied in comparison to the first methodology. However, thermodynamic geometry has two major drawbacks. First, it relies on the slow-driving approximation, which makes it suboptimal for finite-time optimal driving processes, where the driving time is small and the slow-driving assumption is inherently violated. See Ref.[30] for the experimentally verified violation. Second, it lacks the ‘kinks’ in optimal driving protocols, rendering it inconsistent with exact analytical solutions in Ref. [7, 8, 9, 12, 10, 11].
@@
Recently, a third optimal control methodology has emerged in ST, which uses a mathematical framework: the optimal mass transport theory (OMTT) [31, 32, 33, 34]. The insights from OMTT have been incorporated into ST [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. The mapping between ST and OMTT relies on the equivalence between EPR in ST and the Wasserstein distance in OMTT (an information-theoretical distance measure between probability distributions). The optimal transport map corresponds to optimally changing an initial probability distribution to a final probability distribution; equivalently, the optimal driving protocol is obtained here. Hence, the optimal control in OMTT assumes ‘full’ control of the probability distribution defined in the state-space; in comparison, thermodynamic geometry assumes a parametric control defined in the control parameter space. Due to its inherent formulation as an optimization problem, OMTT has a broad range of statistical machine learning applications and advantages [47], to name a few, computer vision, linguistic, signal, and image representation.
@@
Despite its multitude of successful applications, OMTT has three major drawbacks. First, despite the numerical applicability of OMTT, exact analytical solutions are not available except for Gaussian systems, a statistical approximation that does not necessarily hold for finite-size systems in ST prone to non-Gaussian fluctuations [48, 49]. Second, the quadratic dependence of EP on driving, it is an assumption that holds for systems close-to-equilibrium (cEQ) [48, 49, 5, 50], and leads to a massive underestimation of EP for far-from-equilibrium (fEQ) systems [48, 49]. Third, when the driving time to reach the initial to final state is finite, driving time is a resource to be optimized over, OMTT does not take this constraint into account, which results in its suboptimal performance for a finite-time optimal process, since OMTT is built upon on the slow driving assumption. Therefore, the results obtained using OMTT are inconsistent with the model-specific exact analytical results from Ref.[7, 9], the ‘kinks’ are absent even for the most simple models; see Ref.[51] for a comparison.
@@
Due to their better computational aspects, variational formulations of physical processes have been utilized to study the dynamics of non-equilibrium systems, for instance, phase transitions, first-passage times, and metastability in stochastic systems [52, 53, 54, 55, 56, 57, 58, 59]. Variational formulations have broad applicability for efficient numerical optimization problems in machine learning [60, 61, 62], and have recently been explored in ST [63, 64, 65, 66, 67, 68]. Recently, an exact variational formulation for fEQ discrete-state processes has been formulated, namely, the ‘Minimum action principle’ (MinAP) [48, 49], an exact canonical ensemble analog of non-equilibrium discrete-state systems, that connects the transition probability measure to thermodynamic and dynamic physical quantities, namely entropy production and thermodynamic length of transition currents/traffics. The application of MinAP unifies FR and the non-quadratic formulation of TKUR within a single framework [48, 49].
@@
Here, based on MinAP [48, 49], we formulate the generalized finite-time optimal control (GFTOC) framework valid for any discrete-state fEQ system, by exploiting the Riemannian geometric structure in the control parameter space. We compute the optimal finite-time driving protocols that minimize the total driving EP, which exhibit the ‘kinks’. GFTOC unifies the optimal control of slowly driven and finite-time processes, as we prove an exact mapping between them. Importantly, since MinAP assigns a thermodynamic cost for any transition path realization, a thermodynamic cost for sustaining the ‘kinks’ is quantified (a boundary term in driving time), namely ‘thermodynamic shock’. Thereby, the ‘kinks’ reduce the thermodynamic cost of driving (a bulk term in driving time), since the ‘kinks’ reduce the distance to be travelled in the control parameter space. The thermodynamic interplay between the boundary and the bulk terms of driving EP gives rise to the formation of the ‘kinks’ as a dynamic consequence. The dual manifestation of the bulk-boundary term of driving EP has a physical interpretation analogous to the work-heat, but formulated here for the finite-time optimal process in ST. We show that the dynamic counterpart of this thermodynamic phenomena is dictated by the restoration of the ‘finite-time speed limit’: an inherent physical timescale associated with the optimal driving process.
We exhibit the applicability of GFTOC using different examples. We address consistency issues associated with the aforementioned different model-specific formulations of optimal control methodologies, and a unified and universal theoretical framework of optimal control in stochastic thermodynamics is unveiled. Our framework formulates a threefold realization of MinAP in ST as; non-quadratic TKUR, FR, and GFTOC, which correspond to thermodynamic inference, partial control, and full control descriptions, respectively [48]. A ‘full control’ description using GFTOC propels the thermodynamically efficient design and control (in a finite-driving time) of finite-size stochastic systems where stochastic thermodynamics has been experimentally proven theoretical framework. The stochastic thermodynamic insight of the ‘kinks’ formulated by GFTOC develops state-of-the-art of OMTT and Thermodynamic geometry, both of which have exhibited a multitude of practical applications but are based on the slow driving assumption. It is an interesting avenue to explore the practical and interdisciplinary implications of GFTOC in OMTT and Thermodynamic geometry.
Setup
@@
These are quasi-static work, excess, and housekeeping EPR, respectively. They physically correspond to the driving of the functional energy $E$ through a set of external control parameters $\{\lambda_{E}\}$, the relaxation toward the Boltzmann distribution dictated by $E$, and the sustained dissipative transition currents due to non-conservative forces $\{F_{\gamma}\}$, respectively. $\dot{\psi}_{E}$ and $\dot{\Sigma}_{E}^{ex}$ are the boundary terms in the control parameter space and the state space, respectively [5, 49]. $\dot{\psi}_{E}$ is integrated over time to obtain the free-energy difference between the initial and final control parameters, $\psi_{E}(\lambda_{E}^{inl})-\psi_{E}(\lambda_{E}^{fnl})$. Similarly, the KL divergence ($D_{E}^{KL}$) quantifies the statistical distance of the instantaneous state-space distribution from the equilibrium Boltzmann distribution, $\rho_{i}^{E}=e^{-E_{i}+\psi_{E}}$. Thus, a time-integrated excess EPR yields an excess EP, which is the difference between the initial and final state-space distributions and is defined with respect to the initial and final control parameters, $\Sigma_{E}^{ex}=-\Delta_{0}^{\tau}D_{E}^{KL}=D_{E}^{KL}(\lambda_{E}^{inl},t=0)-D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)$, with $D_{E}^{KL}(\lambda_{E}^{inl},t=0)=\sum_{\{\rho_{i}\}}\rho_{i}^{inl}\log{\left(\rho_{i}^{inl}/\rho_{i}^{E}(\lambda_{E}^{inl})\right)}$ and $D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)=\sum_{\{\rho_{i}\}}\rho_{i}^{fnl}\log{\left(\rho_{i}^{fnl}/\rho_{i}^{E}(\lambda_{E}^{fnl})\right)}$. The mean housekeeping EPR $\langle\dot{\Sigma}^{hk}\rangle$ is a non-quadratic function of the non-conservative driving force $F_{\gamma}$ multiplied by the equilibrium traffic $T_{\gamma}^{\perp}$ (the direction orthogonal to the external driving). $T_{\gamma}^{\perp}$ quantifies the scaled equilibrium diffusion constant defined for $\gamma^{\rightleftharpoons}$ due to the equilibrium thermodynamic activity of the states $\rho_{i}$ and $\rho_{j}$. For example, for MJP/ideal CRN, $T_{\gamma}^{\perp}=\rho_{i}+\rho_{j}$, and for interacting CRN, $T_{\gamma}^{\perp}=\rho_{i}e^{\mu_{i}^{int}}+\rho_{j}e^{\mu_{j}^{int}}$, where $\mu_{i}^{int}$ is the chemical potential of $\rho_{i}$ attributed to non-ideal interactions of $\rho_{i}$ [5, 50].
@@
$\begin{split}\mathcal{P}\left[\left\{A_{\gamma},D_{\gamma}\right\}\right]=e^{-\mathcal{S}\left[\left\{A_{\gamma},D_{\gamma}\right\}\right]}.\end{split}$ (3)
where the action $\mathcal{S}\left[\left\{A_{\gamma},D_{\gamma}\right\}\right]=\int_{0}^{\tau}dt\>\mathcal{L}^{*}[\{A_{\gamma},D_{\gamma}\}]$ is a time-integral of the effective transition Lagrangian, which equivalently quantifies the effective mean EPR,

### v3

@@ Abstract
Optimal processes in stochastic thermodynamics are a frontier for understanding the control and design of non-equilibrium systems, with broad practical applications in biology, chemistry, and nanoscale/mesoscale systems. Optimal transport theory and thermodynamic geometry have emerged as optimal control methodology, but they are based on slow-driving and close to equilibrium assumptions. An optimal control framework in stochastic thermodynamics for finite time driving is still elusive. Therefore, we solve in this paper an optimal control problem for changing the control parameters of a discrete-state far-from-equilibrium process from an initial to a final value in finite-time. Optimal driving protocols are derived that minimize the total finite-time dissipation cost for the driving process. Our framework reveals that discontinuous endpoint jumps are a generic, model-independent physical mechanism that minimizes the optimal driving entropy production, whose importance is further amplified for far-from-equilibrium systems. The thermodynamic and dynamic physical interpretation and understanding of discontinuous endpoint jumps is formulated. An exact mapping between the finite-time to slow driving optimal control formulation is elucidated, developing the state-of-the-art of optimal transport theory and thermodynamic geometry, which has been the current paradigm for studying optimal processes in stochastic thermodynamics that relies on slow driving assumptions. Our framework opens up a plethora of applications to the thermodynamically efficient control of a far-from-equilibrium system in finite-time, which opens up a way to their efficient design principles.
@@ 1.1 Background and Motivation
The framework of Stochastic Thermodynamics (ST) has developed the thermodynamic understanding of microscopic/mesoscopic systems [1, 2, 3, 4, 5], thermodynamically-consistent coarse-grained macroscopic systems [5, 6], and the trade-off between thermodynamic length, dissipation, and fluctuations [7, 8]. In ST, the entropy production rate (EPR) quantifies the thermodynamic dissipation cost and constrains the dynamics of non-equilibrium systems. For instance, the fluctuation relation (FR) and the thermodynamic kinetic uncertainty relation (TKUR) have revealed fundamental thermodynamic laws, valid beyond the second law of thermodynamics [1, 2, 3, 4, 5, 9, 7, 8]. Recently, a new paradigm of optimal control problems has emerged in ST. Here, the optimal control problem is loosely defined as transforming an initial state (control parameter) into a final state (control parameter) in a finite time, to compute the optimal driving protocols (change of state/control parameters) that minimize the thermodynamic dissipation. Three classes of methodologies have emerged to study optimal control.
@@
First, optimal processes that minimize finite-time dissipation have been investigated for a stochastic particle in a harmonic trap [10]. Surprisingly, the finite-time optimal protocols were found to exhibit discontinuities at the initial and final times, namely the ‘kinks’. Recently, this mechanism has been understood as a mathematical artefact of the imposed boundary condition [11], and was inspired by model and constraint-specific methodologies developed for quantum systems [12], that are valid under certain approximations, and do not hold for generic cases  [^1] These different methodologies are related to different optimal control formulations of classical stochastic systems discussed subsequently, but they are not necessarily the same. Fundamental differences between them can be drastic, despite their case-specific applicability. First, their specific aims differ significantly depending on what they are designed for, mainly the dynamical aspects; see Ref. [12]. Second, the non-equilibrium notion of entropy production rate itself is not necessarily well defined for quantum thermodynamic systems, but it is of paramount importance for classical stochastic thermodynamic systems. As a result, these quantum methodologies strictly focus on the optimization of dynamical aspects, such as constraints on the dynamical evolution of the quantum state or operators, or on their eigenstates.   These methods justify the importance of dynamical aspects over thermodynamic aspects for finite-time optimal processes, arguing that dynamics become more significant than thermodynamics. However, by construction, the most fundamental principles of statistical physics imply that thermodynamics constrains the dynamics, and not the other way around. Therefore, even the dynamics of finite-time optimal processes ‘must’ be constrained by thermodynamics. We rigorously prove this within our framework. . However, these works fail to capture the physical and thermodynamic origin of the ‘kinks[10, 11]. The universality of ‘kinks’ have been observed in some computationally solvable models [14, 15, 16, 17]. Despite a few model-specific studies, analytically computable solutions for other models are lacking, which has created a void in the theoretical physical understanding/framework to attribute this phenomenon.
@@
Second, thermodynamic length defines the distance in the control parameter space of a model and connects it to the thermodynamic dissipation cost [18, 19, 20, 21, 22, 23]. This Riemannian geometric structure in the control parameter space has been exploited to compute optimal driving protocols and formulate the framework of thermodynamic geometry, which is valid in the slow-driving limit. Within this framework, a geodesic is a minimum-distance path between the initial and final control parameters and is equivalently the optimal driving protocol [23, 24]. Thermodynamic geometry has advantages because of its practical applicability. In particular, metric tensors are numerically/experimentally computed for sophisticated models and systems [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 34]. Therefore, because of its practical applicability, thermodynamic geometry has been rigorously studied in comparison to the first methodology. However, thermodynamic geometry has two major drawbacks. First, it relies on the slow-driving approximation, which makes it suboptimal for finite-time optimal driving processes, where the driving time is small and the slow-driving assumption is inherently violated. See Ref.[37] for the experimentally verified violation. Second, it lacks the ‘kinks’ in optimal driving protocols, rendering it inconsistent with exact analytical solutions in Ref. [10, 11, 14, 17, 15, 16]. Thermodynamic geometry is a sub-case of the general mathematical framework known as information geometry; however, even information geometry does not address these drawbacks [38].
@@
Recently, a third optimal control methodology has emerged in ST, which uses a mathematical framework: optimal transport theory [39, 40, 41, 42]. The insights from optimal transport theory have been incorporated into ST [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. The mapping between ST and optimal transport theory relies on the equivalence between EPR in ST and the Wasserstein distance in optimal transport theory (an information-theoretical distance measure between probability distributions). The optimal transport map corresponds to optimally changing an initial probability distribution to a final probability distribution; equivalently, the optimal driving protocol is obtained here. Hence, the optimal control in optimal transport theory assumes ‘full’ control of the probability distribution defined in the state-space; in comparison, thermodynamic geometry assumes a parametric control defined in the control parameter space. Due to its inherent formulation as an optimization problem, optimal transport theory has a broad range of applications statistical machine learning and computer science [57], to name a few, computer vision, linguistic, signal, and image representation. It is also used to solve nonlinear differential equations [42], which are of paramount importance in physics [39], biology, economics, and fluid mechanics [40].
@@
Despite its multitude of successful applications, optimal transport theory has three major drawbacks. First, despite the numerical applicability of optimal transport theory, exact analytical solutions are not available except for Gaussian systems, a statistical approximation that does not necessarily hold for finite-size systems in ST prone to non-Gaussian fluctuations [7, 8]. Second, the quadratic dependence of EP on the transition affinity, it is an assumption that holds for systems close-to-equilibrium (cEQ) [7, 8, 5, 6], and leads to a massive underestimation of EP for far-from-equilibrium (fEQ) systems [7, 8]. Third, when the driving time to reach the initial to final state is finite, driving time is a resource to be optimized over, optimal transport theory does not take this constraint into account, which results in its suboptimal performance for a finite-time optimal process, since optimal transport theory is built upon on the slow driving assumption. Therefore, the results obtained using optimal transport theory are inconsistent with the model-specific exact analytical results from Ref.[10, 14], the ‘kinks’ are absent even for the most simple models; see Ref.[58] for a comparison.
@@
Due to their better computational aspects, variational formulations of physical processes have been utilized to study the dynamics of non-equilibrium systems, for instance, phase transitions, first-passage times, and metastability in stochastic systems [59, 60, 61, 62, 63, 64, 65, 66], efficient numerical optimization problems in machine learning [67, 68, 69], and recently explored in ST [70, 71, 72, 73, 74, 75]. Recently, an exact variational formulation for fEQ discrete-state processes has been formulated, namely, the ‘Minimum action principle’ (MinAP) [7, 8], an exact canonical ensemble analog of non-equilibrium discrete-state systems, that unifies FR and the non-quadratic formulation of TKUR within a single framework [7, 8].
1.2 Summary of main results
A Overview
@@
Here, based on MinAP [7, 8], we formulate the generalized finite-time optimal control (GFTOC) framework, valid for any discrete-state fEQ system, by exploiting the Riemannian geometric structure in the control-parameter space. We compute the optimal finite-time driving protocols that minimize the total driving EP, which exhibit the ‘kinks’. GFTOC unifies the optimal control of slowly driven and finite-time processes, as we prove an exact geometric mapping between them. Importantly, since MinAP assigns a thermodynamic cost for any transition-path realization, a thermodynamic cost for sustaining the ‘kinks’ is quantified (a boundary term in driving time), namely the thermodynamic shock. Thereby, the ‘kinks’ reduce the thermodynamic cost of driving (a bulk term in driving time), a ‘geometric shortcut’, since the ‘kinks’ reduce the distance to be travelled in the control-parameter space. The thermodynamic interplay between the boundary and the bulk terms of driving EP gives rise to the formation of the ‘kinks’ as a dynamic consequence, and the restoration of ‘the finite-time geometric speed limit’, an inherent physical timescale associated with the optimal driving process, ‘a time-dissipation-geometry tradeoff relation’. The dual manifestation of the bulk–boundary terms of driving EP has a physical interpretation analogous to work–heat, but formulated here for the finite-time optimal driven process in ST.
We demonstrate the applicability of GFTOC through several examples and address consistency issues arising from existing model-specific optimal control formulations. This leads to a unified and universal theoretical framework for optimal control in ST. Our approach provides a threefold realization of the minimum average entropy production (MinAP) principle in ST via the non-quadratic TKUR, fluctuation relations (FR), and GFTOC, corresponding to thermodynamic inference, partial control, and full control descriptions, respectively [7]. The full-control description enabled by GFTOC allows for the thermodynamically efficient design and control of finite-size stochastic systems in finite driving time, where ST has been experimentally validated. The stochastic-thermodynamic insights of the ‘kinks’ in GFTOC advance optimal transport theory and thermodynamic geometry beyond the slow-driving and close-to-equilibrium assumption. Since information geometry accounts only for excess EP and lacks the possibility of kinks [38], GFTOC further advances the state-of-the-art in information geometry.
B Technical outline
We explicitly define five shortcomings of the existing cEQ optimal control methodologies implemented in ST so far.
  • (S1):
    Violation of the Riemannian geometric structure in control parameter space.
  • (S2):
    Exclusion of the housekeeping EPR.
  • (S3):
    Use of quadratic cEQ dissipation functions for the EPR.
  • (S4):
    Constraint of finite driving time.
  • (S5):
    Lack of thermodynamic and dynamic origin and interpretation of kinks.
The exact model-specific results in Ref.[10] focus on resolving $(\text{S}4)$ but do not address $(\text{S}1),(\text{S}2)$ and $(\text{S}5)$. Despite addressing $(\text{S}2)$, optimal transport theory is prone to $(\text{S}1)$, $(\text{S}3)$, and $(\text{S}4)$ 222optimal transport theory introduces an incremental development $(\text{S}3)$ that addresses $(\text{S2})$, however, it is still a fundamental shortcoming for fEQ systems, see Refs.[7, 8]. Therefore, by construction, methodologies with $(\text{S}2)$ are also prone to $(\text{S}3)$. . thermodynamic geometry aims to address $(\text{S}1)$, but remain susceptible to $(\text{S}2)$ and $(\text{S}4)$  [^3] By construction, methodologies with $(\text{S}4)$ are also prone to $(\text{S}5)$..
The key results obtained with our framework are:
  • Section˜2 deals with shortcomings $(\text{S}1),(\text{S}2)$ and $(\text{S}3)$ under the slow-driving assumption. The class of optimal fEQ driving protocols given by universal geodesics (eq.˜10) is defined as the linear interpolation in the geodesic space of the control parameters (eq.˜8). The corresponding dynamic, thermodynamic, and timescale quantities (eq.˜8, eq.˜9, and eq.˜15) are defined as optimal driving protocol, optimal driving EP and geometric speed limit, respectively, quantifying the minimal driving distance path, dissipation, and time.
  • Section˜3 addresses shortcomings $(\text{S}4)$ and $(\text{S}5)$ via an exact mapping to the optimal slow-driving control problem in section˜2. The resulting finite-time dynamical, thermodynamic, and timescale counterparts are given by eq.˜21, eq.˜25, and eq.˜27. The thermodynamic cost (eq.˜24) admits a boundary-bulk decomposition, analogous to heat and work, which is physically associated with the generation of kinks and slow driving, respectively. Thereby, restoring the finite-time geometric speed limit (eq.˜27): ‘a time-dissipation-geometry tradeoff relation’.
  • Section˜4 generalizes the analytical framework to two practically applicable settings: multiple control parameters and novel numerical algorithms. Section˜5 applies the framework to case-specific problems, highlighting novel results and solving several previously open problems.
1.3 Setup
@@
These are quasi-static work, excess, and housekeeping EPR, respectively. They physically correspond to the driving of the functional energy $E$ through a set of external control parameters $\{\lambda_{E}\}$, the relaxation toward the Boltzmann distribution dictated by $E$, and the sustained dissipative transition currents due to non-conservative forces $\{F_{\gamma}\}$, respectively  [^4] Note that, the choice of energy functional $E$ is more flexible and corresponds to the ‘gauge fixing’ and is usually dictated by the experimental/physical constraint. Therefore, it does not necessarily have to be an equilibrium energy functional, for instance, a non-equilibrium analogue of energy is used for a non-equilibrium steady-state $E=ss$, a Lyapunov functional or rate functional that generates the steady-state distribution. This threefold orthogonal decomposition of mean dissipation has been highlighted in Ref.[8], and was generalized to a fourfold orthogonal decomposition of mean dissipation in Ref.[5] to systems that break the action-reaction symmetry.. $\dot{\psi}_{E}$ and $\dot{\Sigma}_{E}^{ex}$ are the boundary terms in the control parameter space and the state space, respectively [5, 8]. $\dot{\psi}_{E}$ is integrated over time to obtain the free-energy difference between the initial and final control parameters, $\psi_{E}(\lambda_{E}^{inl})-\psi_{E}(\lambda_{E}^{fnl})$. Similarly, the KL divergence ($D_{E}^{KL}$) quantifies the statistical distance of the instantaneous state-space distribution from the equilibrium Boltzmann distribution, $\rho_{i}^{E}=e^{-E_{i}+\psi_{E}}$. Thus, a time-integrated excess EPR yields an excess EP, which is the difference between the initial and final state-space distributions and is defined with respect to the initial and final control parameters, $\Sigma_{E}^{ex}=-\Delta_{0}^{\tau}D_{E}^{KL}=D_{E}^{KL}(\lambda_{E}^{inl},t=0)-D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)$, with $D_{E}^{KL}(\lambda_{E}^{inl},t=0)=\sum_{\{\rho_{i}\}}\rho_{i}^{inl}\log{\left(\rho_{i}^{inl}/\rho_{i}^{E}(\lambda_{E}^{inl})\right)}$ and $D_{E}^{KL}(\lambda_{E}^{fnl},t=\tau)=\sum_{\{\rho_{i}\}}\rho_{i}^{fnl}\log{\left(\rho_{i}^{fnl}/\rho_{i}^{E}(\lambda_{E}^{fnl})\right)}$. The mean housekeeping EPR $\langle\dot{\Sigma}^{hk}\rangle$ is a non-quadratic function of the non-conservative driving force $F_{\gamma}$ multiplied by the equilibrium traffic $T_{\gamma}^{\perp}$ (the direction orthogonal to the external driving). $T_{\gamma}^{\perp}$ quantifies the scaled equilibrium diffusion constant defined for $\gamma^{\rightleftharpoons}$ due to the equilibrium thermodynamic activity of the states $\rho_{i}$ and $\rho_{j}$. For example, for MJP/ideal CRN, $T_{\gamma}^{\perp}=\rho_{i}+\rho_{j}$, and for interacting CRN, $T_{\gamma}^{\perp}=\rho_{i}e^{\mu_{i}^{int}}+\rho_{j}e^{\mu_{j}^{int}}$, where $\mu_{i}^{int}$ is the chemical potential of $\rho_{i}$ attributed to non-ideal interactions of $\rho_{i}$ [5, 6].
@@
$\begin{split}\mathcal{P}\left[\left\{A_{\gamma},D_{\gamma}\right\}\right]=e^{-\mathcal{S}\left[\left\{A_{\gamma},D_{\gamma}\right\}\right]},\end{split}$ (3)
where, the action $\mathcal{S}\left[\left\{A_{\gamma},D_{\gamma}\right\}\right]=\int_{0}^{\tau}dt\>\mathcal{L}^{*}[\{A_{\gamma},D_{\gamma}\}]$ is a time-integral of the effective transition Lagrangian, which equivalently quantifies the effective mean EPR,